Barbara Bailey does not believe in analyst predictions. She feels that the company could maintain a constant annual growth rate in dividends per share of 6% in the future or possibly 8% for the next two years and 6% thereafter. Bailey is basing her estimates on an established long-term investment plan into the Latin American, European and Canadian Markets. By venturing into these markets the risk of the firm will increase. The company estimates that the risk premium will increase from 8.9% to 10%. Currently, the company has a risk premium of 6%.

In preparing the long-term financial plan, Bailey Clothiers' chief financial officer has assigned financial analyst, Scott Markham, to evaluate the firm's current stock price. He has asked Scott to consider the conservative predictions of the securities analysts and the aggressive predictions of Barbara Bailey.

Scott has compiled the following information from the 20X9 data to assist in his analysis:

Data: 20X9 Value

Earnings per share (E.P.S.) $ 5.25

Price per share of common stock 42.00

Book value of common stock equity $ 48,000,000

Total common shares outstanding 3,000,000

Common stock dividend per share $ 3.50

  1. Assuming that the securities analysts are correct in predicting that there is no growth in future dividends, what will be the value per share on Bailey Clothiers' stock? (Note: Please use the new required return on the company's stock here.)
  2. If Barbara Bailey's predictions are correct, what will be the value per share of Bailey Clothiers' stock if the firm maintains a constant annual 6% growth rate in future dividends? (Note: Please use the new required return to answer this question.
  3. If Barbara Bailey's predictions are correct, what will be the value per share of Bailey Clothiers' stock if the firm maintains a constant annual growth rate in dividends per share over the next 2 years and 6% thereafter?
Answer #1

a). No-growth dividend model:

Value per share = annual dividend/required return = 3.50/10% = 35

Note: There appears to be a typo in the question as two different risk premiums are mentioned. One as 6%, the other where it is said that risk premium increases from 8.9% to 10%. I am assuming that 10% is the required return as risk-free rate and beta are not provided for calculating the required return (in case 10% is indeed, the risk premium).

b). Constant growth dividend model:

Value per share = current dividend*(1+growth rate)/(required return - growth rate) = 3.5*(1+6%)/(10%-6%) = 61.83

c). Two stage dividend growth model:

Value per share =96.22